Mary's mother is 4 times the age of Mary. The difference of the ages are 28 how old are they?
Answer by stanbon(75887)

Mary's mother is 4 times the age of Mary. The difference of the ages are 28 how old are they?
-------
m = 4M
m - M = 28
---------
Substitute for "m" and solve for "M":
4M-M = 28
3M = 28
M = 9 1/3 years (Mary's age now)
----
Solve for "m":
m = 4M
m = 4(28/3) = 37 1/3 (mother's age now)
--------------------------
Cheers,
Stan H.
==========================